Grade 12th passMagical Mathematics[Interesting Approach]

Find the square roots of 169 by the method of repeated subtraction.

Harshit Singh
5 Years agoGrade 12th pass

1 Answer

Harshit Singh
5 Years ago
Dear Student

169
169–1=168
168–3=165
165–5=160
160–7=153
153–9=144
144–11=133
133–13=120
120–15=105
105–17=88
88–19=69
69–21=48
48–23=25
25–25=0
Here, we have performed subtraction thirteen times.∴√169=13


Thanks